OpenStudy (anonymous):

4x-10≥4-3x A) x≥2 B) x≤2 C) x≥-2 D) x≤-2

OpenStudy (anonymous):

@sadggssa it is similar to one I solved for ya.On same line try this one too,

OpenStudy (anonymous):

\[4x+3x \ge 10+4\]

OpenStudy (anonymous):

\[7x \ge 14\]

OpenStudy (anonymous):

=?

OpenStudy (anonymous):

solve it @sadggssa

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

14/7= 2

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

\[x \ge 14\div7\]

OpenStudy (anonymous):

\[x \ge2\]

OpenStudy (anonymous):

yw! @sadggssa